OpenStudy (anonymous):

What is the 17th term of the arithmetic sequence 18, 22, 26, ...?

OpenStudy (amistre64):

the beneral set up of any arith seq is such that: an = a1 + d(n-1) find the common "d" (difference) and the first term (a1) and you can determine any other term.

OpenStudy (anonymous):

ok so A1 = 18

OpenStudy (anonymous):

because it is the first

OpenStudy (amistre64):

correct; a1 = 18

OpenStudy (amistre64):

what do we have to add to get from a1 to a2? 18 + d = 22

OpenStudy (anonymous):

4

OpenStudy (amistre64):

good, then lets put those into the general set up and see where it takes us an = a1 + d(n-1) an = 18 + 4(n-1) we want the nth term to be the 17th term right? n = 17 so whats a17 turn out to be?

OpenStudy (anonymous):

os so I plugged in 18+4(17-1) and got 22(16) = 352

OpenStudy (anonymous):

But I got it wrong didn't I

OpenStudy (amistre64):

the set up is good, lets just dbl chk the mathing :) a17 = 18 + 4(16) = 18 + 64 = 82

OpenStudy (amistre64):

order of operations is important remember to mulitply or divide first ...

OpenStudy (anonymous):

oh so we plug it into the an ok but how did the 16 come about ... ok

OpenStudy (amistre64):

n-1 is what the general formula states; 17-1 = 16 the nth term in this case is the 17th term

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

youre welcome